# Long-Context Linear System Identification

### v1

@@ 6 Sketch of Proof
$\exists 0<\alpha_{L}<\alpha_{U}\enspace\text{such that}\enspace\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D):\alpha_{L}\leqslant Y_{{\boldsymbol{A}}}\leqslant\alpha_{U}\enspace\text{and}\enspace W_{{\boldsymbol{A}}}\geqslant W_{{\boldsymbol{A}}}^{R}\geqslant W_{{\boldsymbol{A}}}\,.$ (18)
Then, letting $\gamma=2\left(1+2c^{\prime 2}\sigma^{2}\ln dTN\right)$ and $k^{\prime}$ such that $R^{k^{\prime}}\alpha_{L}\geqslant\gamma\alpha_{U}$, we have
$\mathbb{P}\left(W_{\boldsymbol{A}}\leqslant 2\gamma Y_{\boldsymbol{A}}\right)\leqslant\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant\gamma Y_{\boldsymbol{A}}\right)\leqslant\cup_{k=1}^{k^{\prime}}\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant R^{k}\alpha_{L},\gamma Y_{\boldsymbol{A}}\geqslant R^{k-1}\alpha_{L}\right)\,.$
Each of the terms in the union can be controlled by the Freedman’s inequality in Equation 17 with the choices of $r_{Y}=\alpha_{L}R^{k-1}$ and $r_{W}=\alpha_{L}R^{k}$:
$\mathbb{P}\left(W_{\boldsymbol{A}}\leqslant 2Y_{\boldsymbol{A}}\right)\leqslant\sum_{k=1}^{k^{\prime}}\exp\left(-\alpha_{L}R^{k-2}/4\right)\leqslant\exp\left(-\dfrac{\alpha_{L}}{4R}+\ln\ln\dfrac{2e\alpha_{U}}{R\alpha_{L}}\right)\,.$ (19)
As can be seen from Equation 19, the probability of the event {${{\left\{W_{{\boldsymbol{A}}}\leqslant 2Y_{{\boldsymbol{A}}}\right\}}}$ is largely controlled with the lower bound $\alpha_{L}$ as the ratio $\alpha_{U}/\alpha_{L}$ only matters logarithmically. This is crucial as the two bounds differ with the condition number $\kappa$ of the linear operator $L_{\star}$, which can scale with $T$.
Therefore, it possible to control the event $\mathcal{E}$ with a union bound over an $\epsilon$-net of $\mathcal{A}^{\prime}(D)$. In particular, $\alpha_{L}$ needs to be uniformly bounded below such that $\alpha_{L}/R$ is of scale $\ln|\mathcal{N}_{\epsilon}(\mathcal{A}^{\prime}(D))|$. And, this is achieved by Hanson-Wright inequality (Hanson & Wright, 1971) which allows us to derive the needed uniform lower and upper bounds in Equation 18:
$\begin{split}Y_{\boldsymbol{A}}&\leqslant c_{1}\|L_{\star}\|_{\rm op}\cdot\sqrt{pdrN\mathrm{polylog}(p,d,N,T)}{{\left\|\Delta_{{\boldsymbol{A}}}\right\|}}_{F}\,,\\ W_{{\boldsymbol{A}}}^{R}&\geqslant W_{{\boldsymbol{A}}}\geqslant c_{2}\sigma_{\min}(L_{\star})^{2}\cdot N{{\left\|\Delta_{{\boldsymbol{A}}}\right\|}}_{F}^{2}\,,\\ \end{split}$
@@
$\dfrac{\|\Delta_{\boldsymbol{A}}\|_{F}^{2}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}^{2}}\geqslant\ln|\mathcal{N}_{\epsilon}(\mathcal{A}^{\prime}(D))|\,,\quad\text{where}\quad\epsilon\sim\dfrac{\mathrm{polysqrt}(p,d,N,T,e^{\kappa})}{1+c^{2}\ln\frac{1}{\delta}}\,.$
Here, we pick up the dependency on $\kappa$ as $\epsilon$ scales with $\kappa$. This is needed to bound the worst-case errors while transitioning from point-wise bounds on the $\epsilon$-net $\mathcal{N}_{\epsilon}(A^{\prime}(D))$ to the whole set $A^{\prime}(D)$.
@@
$\mathcal{A}^{\prime}(D)={{\left\{{\boldsymbol{A}}\mid\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\geqslant CD^{2}\dfrac{pdr}{N}\mathrm{polylog}(p,d,N,T,\kappa)\right\}}}\,,$
for some $C=\mathcal{O}(\frac{1}{\delta})$ is sufficient to deduce ${\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{A}^{\prime}(D)$ with probability $1-\delta$. The proof of Theorem 4.1 is then complete as
$\|\Delta_{{\boldsymbol{A}}}\|_{F}^{2}=\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{F}^{2}\geqslant(T-p)\|{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}\,.$
7 Conclusion
In this work, we extend non-asymptotic linear system identification theory and derive upper bounds on the sample complexity of learning long-context linear autoregressive models. Our bounds improve upon the existing arguments specific to first-order systems by employing a uniform concentration argument over prediction differences. We further establish improved statistical rates when learning under a low-rank assumption. Finally, we show that even with long or unbounded generative contexts, misspecification still allows the estimation of the matrices with a reduced sample complexity and for stable systems.
While this work makes significant progress for non-asymptotic linear system identification theory, several technical questions remain open for further investigation. Can the OLS operator norm be coarsely controlled to derive rates for unconstrained OLS in the $NT=\Omega(pdr)$ regime? Is it possible to find efficient algorithms that would benefit from low-rank assumptions? Lastly, can misspecification be beneficial for marginally stable systems?
subsubsectionAcknowledgments This project was supported by the Swiss National Science Foundation (grant number 212111).
Appendix A Preliminary tools
A.1 Hanson-Wright inequality
@@
Remark A.2.
For data regime considered in this paper, $K=c\sigma$ in Theorem A.1.
A.2 Freedman’s inequality
We use an extension of Freedman’s inequality (Freedman, 1975) to non-bounded differences by Dzhaparidze & Van Zanten (2001) to show concentration of certain second-order terms. For the sake of completeness, we provide the original Freedman’s inequality. We also remark that it is possible to use the original Freedman’s inequality in our proofs to deal with any bounded noise.
Theorem A.3.
@@ A.2 Freedman’s inequality
Theorem A.4.
(Freedman’s inequality with non-bounded differences) Let $Y_{0},\dots,Y_{n}$ be a real-valued martingale series that is adapted to the filtration $\mathcal{F}_{0},\dots,\mathcal{F}_{n}$ where $Y_{0}=0$. Let $d_{1},\dots,d_{n}$ be the difference sequence induced, i.e.,
@@
$W_{i}^{R}=\sum_{j=1}^{i}\mathbb{E}[d_{j}^{2}\mid\mathcal{F}_{j-1}]+d_{i}^{2}\mathds{1}_{{{\left\{|d_{i}|>R\right\}}}}\quad\text{for }i=1,\dots,n.$ (20)
@@
We extend these theorems in Lemma A.5 to compare the quadratic variation with the martingale series itself. This is useful in our proofs to show certain events necessarily implied by empirical risk minimization do not occur with high probability.
Lemma A.5.
Let $\mathcal{E}$ be the following event
@@
where $10\leqslant\alpha_{L}\leqslant\alpha_{U},R>$ 1 are constants and $W_{n}^{R}$ and $Y_{n}$ are defined as in Theorem A.4. Then, for any $\gamma>0$, we have the following concentration inequality
($\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant\gamma Y_{n}\right\}}}\cap\mathcal{E}\right)\leqslant\exp\left(-\dfrac{\alpha_{L}}{2R\gamma}+\ln\left(\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)+1\right)\right)\,.$
@@
Without loss of generality, we assume $\gamma=1$ as it is possible to scale the martingale series with $\frac{1}{\gamma}$ and set $R\leftarrow\frac{1}{\gamma}R,\alpha_{L}\leftarrow\frac{1}{\gamma^{2}}\alpha_{L}$ and $\alpha_{U}\leftarrow\frac{1}{\gamma}\alpha_{U}$.
Let $\mathcal{G}={{\left\{\alpha_{L},R\alpha_{L},\dots,R^{k}\alpha_{L}\right\}}}$ where $k$ is the smallest positive integer such that
$R^{k}\alpha_{L}\geqslant\alpha_{U}\,.$
@@
$\begin{split}\mathbb{P}\left(W_{n}^{R}\leqslant Y_{n}\cap\mathcal{E}\right)&\leqslant\mathbb{P}\left(\cup_{i=1}^{k}\left({{\left\{W_{n}^{R}\leqslant R^{i}\alpha_{L},Y_{n}\geqslant R^{i-1}\alpha_{L}\right\}}}\cap\mathcal{E}\right)\right)\\ &\leqslant\mathbb{P}\left(\cup_{i=1}^{k}\left({{\left\{W_{n}^{R}\leqslant R^{i}\alpha_{L},Y_{n}\geqslant R^{i-1}\alpha_{L}\right\}}}\right)\right)\\ &\leqslant\sum_{i=1}^{k}\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant R^{i}\alpha_{L},Y_{n}\geqslant R^{i-1}\alpha_{L}\right\}}}\right)\,.\end{split}$
By applying Theorem A.4 with $r=R^{i-1}\alpha_{L}$ and $W=R^{i}\alpha_{L}$, we obtain
$\begin{split}\mathbb{P}\left(W_{n}^{R}\leqslant R^{i}\alpha_{L},Y_{n}\geqslant R^{i-1}\alpha_{L}\right)\leqslant\exp\left(-\alpha_{L}R^{i-2}/4\right)\,,\end{split}$
for each $i=1,\dots,k$. The result follows by noting that
$\begin{split}\sum_{i=1}^{k}\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant R^{i}\alpha_{L},Y_{n}\geqslant R^{i-1}\alpha_{L}\right\}}}\right)&\leqslant\sum_{i=1}^{k}\exp\left(-\alpha_{L}R^{i-2}/4\right)\\ &\leqslant\exp\left(-\dfrac{\alpha_{L}}{4R}+\ln k\right)\\ &\leqslant\exp\left(-\dfrac{\alpha_{L}}{4R}+\ln\left(\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)+1\right)\right)\,.\end{split}$
A.3 Supremum of the noise
@@
Lemma A.6.
Let $X_{1},\dots,X_{n}$ be i.i.d. mean zero and $\sigma^{2}$-sub-Gaussian random variables (in the sense provided in Assumption 3.1). Then, there exist a universal constant $c^{\prime}$ such that for any $t>0$,
$\mathbb{P}\left(\sup_{i=1,\dots,n}X_{i}>c^{\prime}\sigma\sqrt{2\ln n}+t\right)\leqslant 2\exp\left(-\dfrac{t^{2}}{2c^{\prime}\sigma^{2}}\right)\,.$
@@
The result follows by setting $r=c^{\prime}\sigma\sqrt{2\ln n}+t$. ∎
Corollary A.7.
For any $\delta>0$, there exists a universal constant $c^{\prime}(\delta)=\mathcal{O}\left(\sqrt{\ln\frac{1}{\delta}}\right)$ such that
$\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}\leqslant c^{\prime}(\delta)\sigma\sqrt{2\ln dTN}\,,$
@@
Each component $\left(\xi_{t}^{(n)}\right)_{i}$ are i.i.d. of each other and sub-Gaussian with parameter $\sigma$. Therefore, by Lemma A.6, we have
$\mathbb{P}\left(\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}>c^{\prime}\sigma\sqrt{2\ln dTN}+r\right)\leqslant dNT\exp\left(-\dfrac{r^{2}}{2c^{\prime 2}\sigma^{2}}\right)\,.$
Select $r=c^{\prime}\sigma\sqrt{2\left(\ln dTN+\ln\frac{1}{\delta}\right)}$ to obtain the desired confidence level of $\delta$. Note that
$r\leqslant c^{\prime}\sigma\sqrt{2}\left(\sqrt{\ln dTN}+\sqrt{\ln\frac{1}{\delta}}\right)\,,$
and the constant $c^{\prime}(\delta)$ need to satisfy
$c^{\prime}(\delta)\leqslant 1+\sqrt{\dfrac{\ln\frac{1}{\delta}}{\ln dTN}}\leqslant 1+\sqrt{\frac{1}{\ln 2}}\sqrt{\ln\frac{1}{\delta}}\,,$
as $dTN>p\geqslant 2$. Thus, $c^{\prime}(\delta)$ can be picked such that it is a universal constant in $\delta$. ∎
A.4 Proof of Proposition 3.5

### v2

@@ Appendix A Sketch of Proof
$\exists 0<\alpha_{L}<\alpha_{U}\enspace\text{such that}\enspace\forall{\boldsymbol{A}}\in\mathcal{A}^{\prime}(D):Y_{{\boldsymbol{A}}}\leqslant\alpha_{U}\enspace\text{and}\enspace\alpha_{L}\leqslant W_{{\boldsymbol{A}}}\leqslant W_{{\boldsymbol{A}}}^{R}\,.$ (17)
Let $\gamma=C_{1}\ln 2dNT$ and set $k^{\prime}$ to be the smallest integer that verifies $e^{k^{\prime}}\alpha_{L}\geqslant 2\gamma\sigma^{2}\alpha_{U}$. Then,
$\mathbb{P}\left(W_{\boldsymbol{A}}\leqslant 2\sigma^{2}Y_{\boldsymbol{A}}\right)\leqslant\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant 2\gamma\sigma^{2}Y_{\boldsymbol{A}}\right)\leqslant\cup_{k=1}^{k^{\prime}}\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant e^{k}\alpha_{L},2\gamma\sigma^{2}Y_{\boldsymbol{A}}\geqslant e^{k-1}\alpha_{L}\right)\,.$
Each of the terms in the union can be controlled by the Freedman’s inequality in Equation 16 with the choices of $r_{Y}=\alpha_{L}e^{k-1}/2\gamma\sigma^{2}$, $r_{W}=\alpha_{L}e^{k}$ and $R=2e\gamma\sigma^{2}$:
$\begin{split}&\mathbb{P}\left(W_{\boldsymbol{A}}\leqslant 2\gamma\sigma^{2}Y_{\boldsymbol{A}}\right)\leqslant\sum_{k=1}^{k^{\prime}}\exp\left(-\dfrac{e^{k-2}\alpha_{L}}{8e\gamma^{2}\sigma^{4}}\right)\\ &\quad\quad\quad\leqslant\exp\left(-\dfrac{\alpha_{L}}{C_{2}\sigma^{4}\left(\ln 2dNT\right)^{2}}+\ln\left(\ln\left(\dfrac{C_{3}\sigma^{2}\ln\left(2dNT\right)\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)\,.\end{split}$ (18)
As can be seen from Equation 18, the probability of the event {${{\left\{W_{{\boldsymbol{A}}}\leqslant 2\sigma^{2}Y_{{\boldsymbol{A}}}\right\}}}$ is largely controlled with the lower bound $\alpha_{L}$ as the ratio $\alpha_{U}/\alpha_{L}$ only matters logarithmically. This is crucial as the $\alpha_{U}$ and $\alpha_{L}$ differ with the condition number $\kappa$, which can scale with $T$.
Therefore, it possible to control the event $\mathcal{E}$ with a union bound over an $\epsilon$-net of $\mathcal{A}^{\prime}(D)$. In particular, $\alpha_{L}$ needs to be uniformly bounded below as follows:
$\alpha_{L}\geqslant\ln|\mathcal{N}_{\epsilon}(\mathcal{A}^{\prime}(D))|\ln(2dNT)^{2}\,.$
This is achieved by Hanson-Wright inequality (Hanson & Wright, 1971) which allows us to derive the needed uniform lower and upper bounds in Equation 17
$\begin{split}Y_{\boldsymbol{A}}&\leqslant c_{1}\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{pdrN}{{\left\|\Delta_{{\boldsymbol{A}}}\right\|}}_{F}\,,\quad W_{{\boldsymbol{A}}}^{R}\geqslant W_{{\boldsymbol{A}}}\geqslant c_{2}\sigma^{4}\sigma_{\min}(L_{\star})^{2}N{{\left\|\Delta_{{\boldsymbol{A}}}\right\|}}_{F}^{2}\,,\\ \end{split}$
@@
$\dfrac{\|\Delta_{\boldsymbol{A}}\|_{F}^{2}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}^{2}}\geqslant\ln|\mathcal{N}_{\epsilon}(\mathcal{A}^{\prime}(D))|\,,\quad\text{where}\quad\epsilon\sim\dfrac{\mathrm{polysqrt}(p,d,N,T,\kappa)}{1+c^{2}\ln\frac{1}{\delta}}\,.$
Here, a dependency on $\kappa$ is introduced as $\epsilon$ scales with $\kappa$. This is needed to bound the worst-case errors while transitioning from point-wise bounds on the $\epsilon$-net $\mathcal{N}_{\epsilon}(A^{\prime}(D))$ to the whole set $A^{\prime}(D)$.
@@
$\mathcal{A}^{\prime}(D)={{\left\{{\boldsymbol{A}}\mid\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\geqslant C(\delta)D^{2}\dfrac{pdr}{N}\mathrm{polylog}(\kappa,p,d,N,T)\right\}}}\,,$
for some constant $C(\delta)$ is sufficient to deduce ${\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{A}^{\prime}(D)$ with probability $1-\delta$. The proof of Theorem 4.1 is then complete as $\|\Delta_{{\boldsymbol{A}}}\|_{F}^{2}=\|M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\|_{F}^{2}\geqslant(T-p)\|{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}\|_{F}^{2}.$
Appendix B Preliminary tools
B.1 Hanson-Wright inequality
@@
Remark B.1.
For data regime considered in this paper, $K=c\sigma$ in Theorem B.1.
B.2 Freedman’s inequality
We use an extension of Freedman’s inequality (Freedman, 1975) to non-bounded differences by Dzhaparidze & Van Zanten (2001) to show concentration of certain second-order terms. For the sake of completeness, we provide the original Freedman’s inequality. We also remark that it is possible to use the original Freedman’s inequality in our proofs to deal with any strictly bounded noise, leading to improved rates, as explained in Remark C.3.
Theorem B.2.
@@ B.2 Freedman’s inequality
Theorem B.3.
(Freedman’s inequality for non-bounded differences) Let $Y_{0},\dots,Y_{n}$ be a real-valued martingale series that is adapted to the filtration $\mathcal{F}_{0},\dots,\mathcal{F}_{n}$ where $Y_{0}=0$. Let $d_{1},\dots,d_{n}$ be the difference sequence induced, i.e.,
@@
$W_{i}^{R}=\sum_{j=1}^{i}\mathbb{E}[d_{j}^{2}\mid\mathcal{F}_{j-1}]+d_{i}^{2}\mathds{1}_{{{\left\{|d_{i}|>R\right\}}}}\quad\text{for }i=1,\dots,n.$
@@
We extend these theorems in Lemma B.1 to compare the quadratic variation with the martingale series itself. This is useful in our proofs to show certain events necessarily implied by empirical risk minimization do not occur with high probability.
Lemma B.1.
Let $\gamma,R>0,\alpha_{U}\geqslant\alpha_{L}>0$ be scalars and let $\mathcal{E}$ denote the following event
@@
where $Y_{n}$ and $W_{n}^{R}$ verify the assumptions of Theorem B.3. Then, we have the following concentration inequality
($\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant\gamma Y_{n}\right\}}}\cap\mathcal{E}\right)\leqslant\exp\left(-\dfrac{\alpha_{L}}{2e\gamma\left(e\gamma+R\right)}+\ln\left(\ln\left(\dfrac{\gamma\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)\,.$
@@
Let $\mathcal{G}={{\left\{\alpha_{L},e\alpha_{L},\dots,e^{k}\alpha_{L}\right\}}}$ where $k$ is the smallest positive integer such that
$e^{k}\alpha_{L}\geqslant\gamma\alpha_{U}\,.$
@@
$\begin{split}\mathbb{P}\left(W_{n}^{R}\leqslant\gamma Y_{n}\cap\mathcal{E}\right)&\leqslant\mathbb{P}\left(\cup_{i=1}^{k}\left({{\left\{W_{n}^{R}\leqslant e^{i}\alpha_{L},\gamma Y_{n}\geqslant e^{i-1}\alpha_{L}\right\}}}\cap\mathcal{E}\right)\right)\\ &\leqslant\mathbb{P}\left(\cup_{i=1}^{k}\left({{\left\{W_{n}^{R}\leqslant e^{i}\alpha_{L},\gamma Y_{n}\geqslant e^{i-1}\alpha_{L}\right\}}}\right)\right)\\ &\leqslant\sum_{i=1}^{k}\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant e^{i}\alpha_{L},\gamma Y_{n}\geqslant e^{i-1}\alpha_{L}\right\}}}\right)\,.\end{split}$
By applying Theorem B.3 with $r=e^{i-1}\alpha_{L}/\gamma$ and $W=e^{i}\alpha_{L}$, we obtain
$\begin{split}\mathbb{P}\left(W_{n}^{R}\leqslant e^{i}\alpha_{L},Y_{n}\geqslant e^{i-1}\alpha_{L}/\gamma\right)\leqslant\exp\left(-\dfrac{e^{i-2}\alpha_{L}}{2\gamma\left(e\gamma+R\right)}\right)\,,\end{split}$
for each $i=1,\dots,k$. The union bound gives
$\begin{split}\sum_{i=1}^{k}\mathbb{P}\left({{\left\{W_{n}^{R}\leqslant e^{i}\alpha_{L},Y_{n}\geqslant e^{i-1}\alpha_{L}\right\}}}\right)&\leqslant\sum_{i=1}^{k}\exp\left(-\dfrac{e^{i-2}\alpha_{L}}{2\gamma\left(e\gamma+R\right)}\right)\\ &\leqslant\exp\left(-\dfrac{\alpha_{L}}{2e\gamma\left(e\gamma+R\right)}+\ln k\right)\,.\end{split}$
The result follows by noting that
$k\leqslant\ln\left(\dfrac{\gamma\alpha_{U}}{\alpha_{L}}\right)+1\,.$
B.3 Supremum of the noise
@@
Lemma B.2.
Let $X_{1},\dots,X_{n}$ be i.i.d. mean zero and $\sigma^{2}$-sub-Gaussian random variables (in the sense provided in Assumption 3.1). Then, there exists a universal constant $c^{\prime}$ such that for any $t>0$,
$\mathbb{P}\left(\sup_{i=1,\dots,n}|X_{i}|>c^{\prime}\sigma\sqrt{2\ln 2n}+t\right)\leqslant\exp\left(-\dfrac{t^{2}}{2c^{\prime}\sigma^{2}}\right)\,.$
@@
Similarly, we have
$\mathbb{P}\left(\sup_{i=1,\dots,n}-X_{i}\geqslant r\right)\leqslant\cup_{i}\mathbb{P}\left(-X_{i}\geqslant r\right)\leqslant n\exp\left(-\dfrac{r^{2}}{2c^{\prime 2}\sigma^{2}}\right)\,.$
The result follows by union bound with $r=c^{\prime}\sigma\sqrt{2\ln 2n}+t$. ∎
Corollary B.1.
For any $0<\delta<e^{-1}$, there exists a universal constant $c^{\prime}$ such that
$\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}\leqslant c^{\prime}\sigma\sqrt{2}\left(\sqrt{\ln 2dTN}+\sqrt{\ln\dfrac{1}{\delta}}\right)\,,$
@@
Each component $\left(\xi_{t}^{(n)}\right)_{i}$ are i.i.d. and sub-Gaussian with parameter $\sigma$. By Lemma B.2,
$\mathbb{P}\left(\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}>c^{\prime}\sigma\sqrt{2\ln 2dTN}+t\right)\leqslant\exp\left(-\dfrac{t^{2}}{2c^{\prime 2}\sigma^{2}}\right)\,,$
for any $t>0$. Select $t=c^{\prime}\sigma\sqrt{2\ln\frac{1}{\delta}}$ to obtain the desired confidence level of $\delta$. ∎
B.4 Covering numbers
We use the following lemma by (Candès & Plan, 2011) to control the covering numbers of the set of matrices:
Lemma B.3.
(Covering number for low-rank matrices) Let $\mathcal{M}_{r}^{n_{1}\times n_{2}}$ denote the set of low-rank matrices of rank $r$ and Frobenius norm bounded by $1$:
$\mathcal{M}_{r}^{n_{1}\times n_{2}}={{\left\{M\in\mathbb{R}^{n_{1}\times n_{2}}:\|M\|_{F}=1,{\rm rank}(M)\leqslant r\right\}}}\,.$
Then, there exists an $\epsilon$-net $\mathcal{S}$ of $\mathcal{M}$ for Frobenius norm such that
$|\mathcal{S}|\leqslant\left(\dfrac{9}{\epsilon}\right)^{(n_{1}+n_{2}+1)r}\,.$
Corollary B.2.
Let $\mathcal{M}_{r,p^{\prime}}(D)$ denote the following set of matrices with bounded operator norm:
$\mathcal{M}_{r,p^{\prime}}(D)={{\left\{{\boldsymbol{A}}\in\mathbb{R}^{d\times pd}\mid\|{\boldsymbol{A}}\|_{\rm op}\leqslant D,\ {\rm rank}(A_{i})\leqslant r,\ A_{p^{\prime}+1}=\cdots=A_{p}=0\right\}}}\,.$
Then, there exists an $\epsilon$-net $\mathcal{S}$ of $\mathcal{M}_{r,p^{\prime}}(D)$ for operator norm such that
$|\mathcal{S}|\leqslant\exp\left(9p^{\prime}dr\ln\dfrac{Dp^{\prime}}{\epsilon}\right)\,.$